### v1

@@ Appendix B Achieving the lower bounds: Preconditioning and initialization
Theorem B.1 (Preconditioned TARK with volume sampling).
@@
The problem of approximately solving a least-squares problem from a small number of entry evaluations of the vector $\boldsymbol{b}$ has also received recent attention in the context of active learning [6, 19]. Existing approaches achieve the guarantee $\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}\leq(1+\varepsilon)\cdot\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ with high probability after accessing just $\mathcal{O}(r/\varepsilon)$ [6] or $\mathcal{O}(r\log r+r/\varepsilon)$ [27, 9] entries of $\boldsymbol{b}$. Compared to this previous work, Theorem B.1 attains nearly the optimal rate and is among the simplest and most explicit bounds for active linear regression methods.
Appendix C Proofs for ridge regression
This section proves the RK-RR and TARK-RR error bounds. The analysis follows a pattern similar to Section 2.1, with two lemmas followed by the proofs of Theorems 3.1 and 3.2. The two lemmas are minor modifications of Lemmas 2.1 and 2.3. However, the proof of Theorem 3.1 requires a new strategy, since there is not a simple one-step recursion bounding $\operatorname{\mathbb{E}}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\mu}\right\|^{2}$ in terms of $\operatorname{\mathbb{E}}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{\mu}\right\|^{2}$. Instead, it is necessary to use a bias–variance decomposition inspired by [7, 14].
Lemma C.1 (Multi-step expectations).
@@ Appendix C Proofs for ridge regression
Lemma C.3 (Demmel condition number bound).
@@
Proof C.5 (Proof of Theorem 3.1).
@@
To bound the square bias term, average over the random index $i_{t}$ and apply Lemma C.3:
@@
The analysis of the variance is more delicate. Since $\boldsymbol{v}_{t}$ follows the same recurrence as $\boldsymbol{x}_{t}$, the relation Eq. 14 from the proof of Lemma C.1 can also be applied to $\boldsymbol{v}_{t}$, yielding
@@
Proof C.6 (Proof of Theorem 3.2).
@@
Next analyze the terms $\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\mu})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\mu})\bigr{]}$ for $r\leq s$ using Lemma C.1 and Theorem 3.1:

### v2

@@ Appendix C Achieving the lower bounds: Preconditioning and initialization
Theorem C.1 (Preconditioned TARK with volume sampling).
@@
The problem of approximately solving a least-squares problem from a small number of entry evaluations of the vector $\boldsymbol{b}$ has also received recent attention in the context of active learning [6, 19]. Existing approaches achieve the guarantee $\lVert\boldsymbol{b}-\boldsymbol{A}\widehat{\boldsymbol{x}}\rVert^{2}\leq(1+\varepsilon)\cdot\lVert\boldsymbol{b}-\boldsymbol{A}\boldsymbol{x_{\star}}\rVert^{2}$ with high probability after accessing just $\mathcal{O}(r/\varepsilon)$ [6] or $\mathcal{O}(r\log r+r/\varepsilon)$ [27, 9] entries of $\boldsymbol{b}$. Compared to this previous work, Theorem C.1 attains nearly the optimal rate and is among the simplest and most explicit bounds for active linear regression methods.
Appendix D Proofs for ridge regression
This section proves the RK-RR and TARK-RR error bounds. The analysis roughly parallels the analysis in Section 2.1. However, the proof of Theorem 3.1 requires a new strategy, since there is not a simple one-step recursion bounding $\operatorname{\mathbb{E}}\left\|\boldsymbol{x}_{t+1}-\boldsymbol{x}_{\mu}\right\|^{2}$ in terms of $\operatorname{\mathbb{E}}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{\mu}\right\|^{2}$. Instead, it is necessary to use a bias–variance decomposition inspired by [7, 14].
Lemma D.1 (Multi-step expectations).
@@ Appendix D Proofs for ridge regression
Lemma D.3 (Demmel condition number bound).
@@
Proof D.5 (Proof of Theorem 3.1).
@@
To bound the square bias term, average over the random index $i_{t}$ and apply Lemma D.3:
@@
The analysis of the variance is more delicate. Since $\boldsymbol{v}_{t}$ follows the same recurrence as $\boldsymbol{x}_{t}$, the relation Eq. 14 from the proof of Lemma D.1 can also be applied to $\boldsymbol{v}_{t}$, yielding
@@
Proof D.6 (Proof of Theorem 3.2).
@@
Next analyze the terms $\operatorname{\mathbb{E}}\bigl{[}(\boldsymbol{x}_{r}-\boldsymbol{x}_{\mu})^{\top}(\boldsymbol{x}_{s}-\boldsymbol{x}_{\mu})\bigr{]}$ for $r\leq s$ using Lemma D.1 and Theorem 3.1: